OpenStudy (anonymous):

My class is working with SSS and SAS, I was wondering if somebody could help me with two of my proofs? They are in the comments section.

OpenStudy (anonymous):

Show that the triangles are congruent for the given value of the variable.

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\triangle GHI \cong \triangle IHJ, x=4\]

OpenStudy (xapproachesinfinity):

hmmm that your given \(\triangle GHI \cong \triangle IHJ\) does not seem right

OpenStudy (anonymous):

No, they want me to prove that.

OpenStudy (xapproachesinfinity):

well it can't be if your diagram is correct

OpenStudy (xapproachesinfinity):

you mean \(\triangle GHJ\cong\triangle IHG\)

OpenStudy (anonymous):

Oh, yeah. Sorry.

OpenStudy (xapproachesinfinity):

Okay so x=4 you say

OpenStudy (anonymous):

Yes, and I need to use that to write a two-column proof showing that the triangles are congruent.

OpenStudy (xapproachesinfinity):

if x=4 then 2x-9=3 and 2x-3=5 that means \(HG\cong HI\) and \(GJ\cong IJ\)

OpenStudy (xapproachesinfinity):

now we also have HJ is a common segment then by SSS the two triangles are congruent

OpenStudy (xapproachesinfinity):

that's the proof

OpenStudy (anonymous):

Ok, got it thank you.

OpenStudy (xapproachesinfinity):

welcome

OpenStudy (anonymous):

@Directrix could you help me with this one. They give x=4 They want me to prove \[\triangle RST \cong \triangle TUR\]

OpenStudy (anonymous):



OpenStudy (anonymous):

My bad looking at the last one x=18

OpenStudy (xapproachesinfinity):

eh it's the same thing now they want you to do SAS

OpenStudy (anonymous):

Yes, they gave me x=18.

OpenStudy (xapproachesinfinity):

4x-11=61 2x=36 what does mean

OpenStudy (xapproachesinfinity):

UT=RS and \(\angle RTU\cong \angle TRS\)

OpenStudy (anonymous):

4(18)-11=61 - RS=UT 2(18)=36 - \[\angle RTU \cong \angle TRS\]

OpenStudy (xapproachesinfinity):

you need one more side to have SAS can you figure it out

OpenStudy (anonymous):

No it is not given to be a parallelogram @Directix, sorry wasn't trying to ignore you, just trying to keep up with both of you. I do not know how to get the other sides, that is why I'm so confused @xapproachesinfinity

OpenStudy (xapproachesinfinity):

hmmm lol @Directrix we can't assume it is a quadrilateral we one to prove that by proving that the two triangle are congruent

OpenStudy (xapproachesinfinity):

well again there is a common side btw the two triangle

OpenStudy (xapproachesinfinity):

so by sas they are indeed congruent

OpenStudy (anonymous):

Oh, gosh I am always over thinking things, I kept ignoring that side and wanted to try proving RU=ST. LOL

OpenStudy (xapproachesinfinity):

hmm... you need to get an easier way always if it is available

Directrix (directrix):

>>> we can't assume it is a quadrilateral we one to prove that by proving that the two triangle are congruent FALSE

OpenStudy (xapproachesinfinity):

this is a two column proof so no need to go further proving a hard side

OpenStudy (anonymous):

Thank you both for helping me, I have to go now. Dinner is ready.

OpenStudy (xapproachesinfinity):

of course we still need more to prove it is quadrilateral our concern here is jut the concurrency of the two triangles

OpenStudy (xapproachesinfinity):

YW!